• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name X. Zheng

    Log entry time 15:59:49 on March24,2007

    Entry number 197276

    keyword=day shift summary

    8:40 we finished kinematics #11 (last run is 1915). Start moving HRSR to kin #6. Meantime asked MCC to do multi harp scan with empty target in beam;
    9:34 HRSR is at 71.44 (floor mark 71.0 and vernier +42); Harp scan is still ongoing;
    9:50 MCC called and said the PMT HV for the superharp has turned itself off. They cannot reset it from there. We looked into it but cannot reset it either. MCC will try to get the data without cycling the crates;
    10:06 Harp scan is finished. Energy spread is the same as before and the beam spot size in y is 150um;
    10:08 move target to Pb3 and start kin#6;

    11:30 MCC will stop the beam and fix the energy spread, it will take about 30 minutes.

    13:32 MCC called and said when the beam trips next time, they will put the OTR in for 10sec to check the beam spot shape. I was told it should not affect our data taking (since it's only 10sec).

    14:20 MCC want to fix the energy spread again, this was between run 1923 and 1924. The energy spread shown on (1C12) screen is now 6.3x10^-5, similar to yesterday evening. We were concerned that this is higher than the number on the whiteboard (<5e-5), but was told this is not the absolute energy spread and it is okay as long as there is no significant change/increase/jump in this number.

    16:00 By the end of shift, we need about 0.45 C more, or 5-6 more runs (starting from run 1926) for kinematics #6.

    Run summary:

    run 1915, 40uA, 1M events, 0.054C

    run 1916, 40uA, 80K events (current ramp-up), 0.01C
    run 1917, 40uA, 1M events, 0.076C
    run 1918, 40uA, 1M events, 0.07C

    run 1919, 10uA, 80K events, first run after fixing energy spread, 0.005C
    run 1920, 20uA, 122K events, 0.01C
    run 1921, 40uA, 1M events, 0.08C
    run 1922, 40uA, 1M events, 0.08C
    run 1923, 40uA, 1M events, 0.08C
    run 1924, 40uA, 1M events, 0.08C
    run 1925, 40uA, 1M events, 0.08C